A salmon fishing boat on a BC river travelled 96 km upstream (against the current in 4h. Returning downstream at the same speed, it took 3h. Find the speed of the fishing boat in still water and the speed of the river's current
Answer by mananth(16949)

Boat speed speed =x mph
current speed speed =y mph
against wind x-y 4.00 hours
With current x+y 3.00 hours

Distance = same= 96 miles
t=d/r
96 / ( x - y )= 4.00
4 ( x - y ) = 96.00
4 x -4 y = 96 ....................1
96 / ( x + y )= 3.00
3.00 ( x + y ) = 96
3.00 x + 3.00 y = 96 ...............2
Multiply (1) by 3.00
Multiply (2) by 4.00
we get
12 x + -12 y = 288
12 x + 12 y = 384
24 x = 672
/ 24
x = 28 mph

plug value of x in (1)
4 x -4 y = 96
112 -4 y = 96
-4 y = 96 -112
-4 y = -16
y = 4 mph